OpenStudy (vicky123):

solve 5x+6/4 + 8x+6/10 = -2

OpenStudy (anonymous):

got this

OpenStudy (anonymous):

5x+6/4 + 8x+6/10 = -2 (5x + 8x) + (6/4+6/10) = -2 13x +(3/2 + 6/10) = -2 13x + (15/10 + 6/10) = -2 13x + 21/10 = -2 13x = -(4 + 1/10) x = -41/130